Circular No. 6114

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1994ah IN ESO 356-G19 R. H. McNaught reports his discovery of an apparent supernova (mag

∼ 18) on a U.K. Schmidt Telescope I plate taken on Oct. 13. SN 1994ah is at α = 2^{h}54^{m}35^{s}.26, δ = −34^{o}06′45′′.8 (equinox 1950.0; uncertainty 0′′.3 in each coordinate), which is 12′′.4 east and 2′′.5 north of the galaxy’s center. No image appears in this position on the ESO B or R survey plates, the Whiteoak extension of the Palomar Sky Survey, the SERC J survey, or a U.K. Schmidt plate taken on 1993 Oct. 21. Confirmation was obtained by N. B. Suntzeff on Dec. 12.2 UT with the Cerro Tololo 4-m telescope; his spectrophotometry (range 320-750 nm, resolution 700) shows this object to be a type-Ia supernova ∼ 4 months past maximum. Its spectrum appears identical to the nebular spectrum of SN 1981B at the same phase of evolution (Branch et al. 1983, Ap.J. 270, 123).

NOTICE TO SUBSCRIBERS Owing to increased postal rates to be introduced next month, it is

necessary to increase the subscription rates for these Circulars. The new monthly rates will be $15.00 (regular or invoiced) and $9.00 (special or non-invoiced). However, the monthly charge for the Computer Service/email delivery will be decreased to $10.00 (regular) and $6.00 (special). This reduction means that there will be no change in the combined cost to those who subscribe to both the printed and e-mail versions. A subscriber who receives only the e-mail version and who has paid ahead is invited to renegotiate the expiry date for his or her account by sending an e-mail message to iausubs@cfa.harvard.edu by 1995 Jan. 27, indicating either the account number or the e-mail address used for the delivery of the Circulars.

Long-term subscribers will note that the cost of a Computer Service/email subscription is as low as it has ever been, and it could be reduced further if more people who read “pirated” versions of these Circulars were to participate. The Computer Service has many features now it did not have when it started in 1984, including a complete set of the Circulars back to 1982 and a complete cross-referenced index back to 1922. E-mail subscribers who do not have instructions for using the Computer Service should request their passwords by e-mailing the aforementioned address.

1994 December 12 Daniel W. E. Green